OpenStudy (anonymous):

I need your help! How exactly would do this? What is the slope of the line passing through the points (2, 7) and (-1, 3)? o 2/7 o 3/4 o 4/3 o 1/3

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

So, it would be: m = 3 - 7 over -1 - 2 m = -4 over -3 It's negative though. How would it turn positive?

OpenStudy (anonymous):

Yup...

OpenStudy (anonymous):

What would make it turn positive?

OpenStudy (anonymous):

\[\frac{-1}{-1}=1\]

OpenStudy (anonymous):

Oh okay, I forgot about that. Thank you very much! :)